GROUND-STATE SPLITTING OF Am^{2+} IN CaF_{2}

\underline{SCOTT R. BROZELL} AND RUSSELL M. PITZER, Department of Chemical Physics, The Ohio State University, Columbus, OH 43210.

The energy-level splittings of the 5f^{7} nominal ^{8}S_{7/2} ground-state of Am^{2+} in an octahedral site of CaF_{2} were studied using ab initio quantum chemical methods. The CaF_{2} host was modeled with a large finite cluster of ions which approximate the Madelung potential of the crystal lattice^{a}. The potential energy curve for the symmetric stretch of the nearest-neighbor fluoride ions was calculated. The actinide dopant was treated with relativistic core potentials and Gaussian double-zeta basis sets.

aN. W. Winter, M. Ross and R. M. Pitzer J. Phys. Chem. 94, 1172 (1990).